a father is thirty times older than his son. However 18 years later he will only thrice as old as the son. What is the fathers present age
Answer by Edwin McCravy(20056)

This is a weird problem.  If the son were only 2, the father would be 60.
60-year old men normally do not have children that young, so to be true to
life, the son can only be a baby.  Ages of babies are given in months, so 
let's work the problem in months.

Son's age in months = x
Father's age in months = 30x
Son's age in 18 years = Son's age in 18*12 or 216 months = x+216
Father's age in 18 years = Father's age in 216 months = 30x+216

30x + 216 = 3(x + 216)
30x + 216 = 3x + 648
      27x = 432
        x = 16

So the child is a small tot 16 months old, or 1 year and 4 months old.

Answer:  The father is 30*16 or 480 months or 480/12 = 40 years old

Checking:

a father is thirty times older than his son.
40 is 480 months old, which is 30 times 16 months.

In 216 months, the child will be 232 months old, or 19 years and 4 months
and the father will be 480+216 or 696 months or 58 years old.

It's right but the problem is weird.

Edwin